### v1

@@ 12. University of Vienna, Department of Astrophysics, Türkenschanzstraße 17, 1180 Wien, Austria
(Received XX XX XXXX / Accepted XX XX XXXX)
We present a reconstruction of the line-of-sight motions of the local interstellar medium (ISM) based on the combination of a state-of-the-art model of the three-dimensional dust density distribution within 1.25 kpc from the Sun and the Hi and CO line emission within Galactic latitudes $|b|$$\leq$ 5. We use the histogram of oriented gradient (HOG) method, a computer vision technique for evaluating the morphological correlation between images, to match the plane-of-the-sky 3D dust distribution across distances with the atomic and molecular line emission. We identify a significant correlation between the 3D dust model and the line emission. We employ this correlation to assign line-of-sight velocities to the dust across density channels and produce a face-on map of the local ISM radial motions with respect to the local standard of rest (LSR). We find that most of the material in the 3D dust model follows the large-scale pattern of Galactic rotation; however, we also report local departures from the rotation pattern with standard deviations of 12.1 and 6.1 km s-1 for the Hi and CO line emission, respectively. The mean kinetic energy densities of these streaming motions are around 0.68 and 0.18 eV/cm3 from either gas tracer. Assuming homogeneity and isotropy in the velocity field, these values are within a factor of a few of the total kinetic energy density. These kinetic energy values are roughly comparable to other energy densities. thus confirming the near-equipartition introduced by the feedback loops connecting the physical processes in the local ISM. Yet, we find energy and momentum overdensities of around a factor of ten concentrated in the Radcliffe Wave, the Split, and other local density structures, particularly around the Vela C and Ara star-forming regions. Although we do not find evidence of the Local Spiral Arm impact in these overdensities, their distribution suggests the influence of other large-scale effects that, in addition to supernova feedback, shape the energy distribution and dynamics in the Solar neighborhood.
@@ 1 Introduction
Tracing the flows of matter and energy in the interstellar medium (ISM) is instrumental for understanding the structure and evolution of the Milky Way and other galaxies (see, for example, Ballesteros-Paredes et al., 2020; Tacconi et al., 2020; Chevance et al., 2023). Reconstructing the flow of gas in the ISM enables the identification of the mechanisms that make gas available for star formation and allows us to quantify the impact of galactic dynamics, magnetic fields, stellar winds, outflows, and supernovae (SNe) in the ISM (e.g., Field & Saslaw, 1965; Tomisaka, 1984; Koyama & Inutsuka, 2000; Blitz et al., 2007; Hennebelle et al., 2008; Klessen & Hennebelle, 2010; Dobbs et al., 2014; Krumholz et al., 2014; Klessen & Glover, 2016; Schmidt et al., 2016; Hennebelle & Inutsuka, 2019). In this work, we present a reconstruction of the ISM line-of-sight (LOS) motions in the Solar neighborhood using state-of-the-art models of the three-dimensional dust density distribution, which we associate with the velocity information from the neutral atomic hydrogen (Hi) and carbon monoxide (CO) line emission using a machine vision tool to quantify the morphological similarities between distance and velocity channels (Soler et al., 2019).
Classic studies of the ISM distribution and kinematics rely on the assumption of circular motions around the Galactic center to compute what is usually called “kinematic distances” (see, for example, Oort et al., 1958; Roman-Duval et al., 2009; Sofue, 2011; Wenger et al., 2018; Hunter et al., 2024). Kinematic distances depend on the Galactic rotation model and are affected by the kinematic distance ambiguity (see, for example, Eilers et al., 2020; Reid, 2022). Moreover, the motion of the Galactic ISM is not purely rotational; non-circular streaming motions are induced by the Galactic bar, the spiral arms, and stellar feedback (Burton, 1971; Gómez, 2006; Moisés et al., 2011). The latter effect is expected to be prominent in the Solar neighborhood, where the motions introduced by supernovae (SNe)-blown bubbles and other local effects are potentially dominant over Galactic rotation as seen from the local standard of rest (LSR, see, for example, Zucker et al., 2023, and references therein).
@@
In this work, we associate the state-of-the-art 3D dust models presented by Edenhofer et al. (2024) with the LOS velocity information provided by Hi and CO surveys toward the Galactic plane. Following the pioneering work in Tchernyshyov & Peek (2017); Tchernyshyov et al. (2018), we adopt the term “kinetic tomography” to describe our reconstruction of the density and LOS velocity distribution. Our approach is based on the morphological similarity quantified by the histogram of oriented gradients (HOG) method (Soler et al., 2019). Using HOG, we link the position-position-distance (PPD) 3D dust density cubes and the position-position-velocity (PPV) Hi and CO line emission cubes. The result is a position-position-distance-velocity (PPDV) hypercube, which we employ to characterize the local ISM motions along the line of sight. Line emission from ISM tracers and 3D dust have previously been combined to investigate the ISM in four dimensions (see, for example, Ivanova et al., 2021; Duchêne et al., 2023), but not with the level of reconstruction in the Edenhofer et al. (2024) maps, and never before with the HOG method.
@@
We use the Hi line emission at 21-cm wavelength, which traces both the cold, pre-molecular state before star formation and the warm, diffuse ISM before and after star formation (see, McClure-Griffiths et al., 2023, for a recent review). Since its discovery, Hi has been instrumental in studying the diffuse ISM in the Milky Way (Ewen & Purcell, 1951; Muller & Oort, 1951; Pawsey, 1951). Early observations of the Hi absorption against radio continuum sources revealed the presence of narrow, few-km s-1-wide, spectral features (e.g., Hagen et al., 1955; Clark, 1965). In emission, these narrow features appear on top of broader 10 to 20-km s-1-wide features (e.g., Matthews, 1957). These observations inspired the formulation of a “two phase” Hi model (Field et al., 1966; McKee & Ostriker, 1977), in which at the pressure of the ISM, the heating and cooling processes naturally lead to two thermally stable states: a dense cold neutral medium (CNM; $T$$\approx$ 50 K and $n$$\approx$ 50 cm-3) immersed in a diffuse warm neutral medium (WNM; $T$$\approx$ 8 000 K and $n$$\approx$ 0.3 cm-3).
@@
Part of our understanding of the Hi multiphase structure comes from studying absorption toward continuum sources. (see, for example, Strasser et al., 2007; Stanimirović et al., 2014; Murray et al., 2018). Comparisons between 21-cm Hi emission and absorption measurements indicate that, in the vicinity of the Sun, the WNM has roughly the same column density as the CNM (Falgarone & Lequeux, 1973; Liszt, 1983). Crucial additional information about the distribution of the CNM in and around MCs comes from the portion of the CNM sampled by the extended absorption of background Hi emission by cold foreground Hi, which is generically known as Hi self-absorption (HISA; Heeschen, 1955; Gibson et al., 2000; Seifried et al., 2020). Observations indicate that the HISA distribution is related to that of the CO emission, although the CNM it reveals corresponds to less than 15% than the total gas mass (see, for example, Gibson et al., 2005; Krčo et al., 2008; Wang et al., 2020b; Syed et al., 2020).
@@
We also use the CO $(J$$=$$1$$\rightarrow$0)0)0 ) emission at 2.6 mm wavelength, which is the archetypal tracer of the cooler and denser ISM (Wilson et al., 1970). With its low excitation energy and critical density, CO is easily excited and provides an irreplaceable proxy for H2, which is the most abundant molecule in the Galaxy but is much harder to observe (Combes, 1991; Bolatto et al., 2013; Heyer & Dame, 2015). The focus on the molecular material traced by CO in star formation (SF) studies was cemented by the observed correlation between the SF and CO surface densities in nearby galaxies, which contrasts with the lack of correlation between SF and Hi at low surface densities in those objects (see, for example, Kennicutt, 1998; Wong & Blitz, 2002; Leroy et al., 2008; Schinnerer & Leroy, 2024).
@@
In this paper, we apply the HOG method to study the atomic and molecular gas motions toward the nearby Milky Way’s disk, defined as the Galactic latitude range $|b|$$\leq$ 5 within the 1.25-kpc extent of the Edenhofer et al. (2024) 3D dust reconstruction. The presentation of the data, analysis methods, and results is organized as follows. In Sec. 2, we present the 3D dust models and the Hi and CO observations used in the analysis. Section 3 summarizes the main aspects of the HOG method implementation for analyzing the Galactic plane. We describe the morphological correlation between the 3D dust models and the line emission in Sec. 4. In Sec. 5, we detail the streaming motions obtained with the HOG method and the energy and momentum densities derived from them. Section 6 discusses our results and their implications for understanding the ISM dynamics in the Solar neighborhood. We present our conclusions of this work in Sec. 7. We complement the main results of this work with the analysis shown in a set of appendices. Appendix A presents details on the HOG method’s error propagation and selection of parameters. We consider the effect of HISA in our results and the potential of the HOG method to identify these features in App. B. In App. C, we describe the assumptions for calculating the physical quantities derived with the HOG-based LOS velocity reconstructions. Appendix D compares our results with the distances and line-of-sight velocities for the five maser sources within the studied volume. Finally, App. E presents the distance-velocity mapping for a few regions of interest, including areas toward the Galactic center and anticenter, where uncertainties in the kinematic distance estimates are considerable.
@@ 2 Data
Throughout this paper, we use $v_{\rm LOS}$ to refer to the radial velocity inferred from the emission line Doppler shift relative to the LSR, usually identified as $v_{\rm LSR}$ in the Hi and CO observations. We introduce this convention to distinguish it from the general motion relative to the LSR, for which the three components of the velocity vector can be inferred in the Solar vicinity from the measurement of stellar motions (see, for example, Miret-Roig et al., 2022; Ratzenböck et al., 2023).
The selected $b$ range implies that we are studying the volume covered by the revolution of a trapezoid with base sizes 12.16 and 220.4 pc located at 69 and 1250 pc from the LSR, as illustrated in Fig. 2. The limits along the LOS are set by the boundaries of the Edenhofer et al. (2024) 3D dust model. At the furthest point in this analysis, we will be covering a region roughly within the characteristic CO scale height around the Solar circle ($\sim$ 100 pc, Heyer & Dame, 2015).
We consider the line emission in the range $-25$$<$$v_{\rm LOS}$$<$$25$ km s-1, which corresponds to the expected amplitude of LOS velocities for a Galactic rotation model for heliocentric distances within 1250 pc plus or minus a few kilometers per second. Although this selection limits the range of the recovered departures from Galactic rotation, it reduces the effect of chance correlation in the morphological matching, as discussed in App. A. We also present the results of broader input $v_{\rm LOS}$ ranges in App A.3.3.
@@
The primary dataset enabling our study is the reconstruction of the 3D dust density between 69 and 1250 pc from the Sun presented by Edenhofer et al. (2024). This model uses the extinction estimates from Zhang et al. (2023), which are primarily based on the Gaia satellite’s BP/RP spectra, with an spectral resolution ($R$) of roughly between 30 and 100 (Gaia Collaboration et al., 2023). Zhang et al. (2023) forward-modeled the extinction, distance, and intrinsic parameters of each star given the combination of the Gaia spectra (Carrasco et al., 2021; De Angeli et al., 2023; Montegriffo et al., 2023) and infrared photometry from the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006), and unWISE, a processed catalog based on the observations from NASA’s Wide-field Infrared Survey Explorer (WISE) mission (Wright et al., 2010; Schlafly et al., 2019). The model was trained using a subset of stars observed with higher spectral resolution, $R$$\sim$ 1800, available from the Large Sky Area Multi-object Fibre Spectroscopic Telescope (LAMOST; Wang et al., 2022; Xiang et al., 2022).
@@
Edenhofer et al. (2024) obtains a spherical-coordinate reconstruction beyond 1 kpc while still resolving nearby dust clouds at parsec-scale resolution by implementing a new Gaussian process (GP) methodology to incorporate smoothness in a spherical coordinate system, mitigating fingers-of-god artifacts. The authors modeled the 3D distribution of differential extinction for stars in the Zhang et al. (2023) catalog, assuming that the dust extinction distribution is spatially smooth. The posterior of their extinction model was reconstructed using variational inference and Gaussian variational inference (MGVI, Knollmüller & Enßlin, 2019; Leike & Enßlin, 2019). We refer to Edenhofer et al. (2024) for further details on the data modeling and reconstruction.
The result of the Edenhofer et al. (2024) model is a set of 12 samples drawn from the variational posterior of the 3D dust extinction distribution; that is, a set of 12 possible distributions consistent with the observations within the uncertainties. Each sample gives the value of the dust density in voxels arranged as a series of HEALPix [^1] HEALPix is the Hierarchical Equal Area isoLatitude Pixelization; see http://healpix.sf.net and Górski et al. (2005). spheres. There are 516 of these spheres, each corresponding to a distance logarithmically-spaced grid between 69 to 1250 pc. Each sphere has a resolution parameter $N_{\rm side}$$=$ 256, corresponding to an angular size of around 13′7 for each voxel. We analyzed each of the 3D dust extinction samples, considering them independent realizations of 3D dust extinction, and reported the mean trends, as described in detail in App. A.
Using the tools in the Python healpy package, we produced a Cartesian projection of the 3D dust extinction covering the range $-180$$<$$l$$<$ 180 and $-5$$<$$b$$<$ 5 with a pixel size $\delta l$$=$$\delta b$$=$ 7.5. The reconstruction was made in terms of the unitless extinction defined by Zhang et al. (2023), which we transformed into V-band extinction ($A_{\rm V}$) by multiplying with the 2.8 factor derived from their extinction curve. We then transformed $A_{\rm V}$ into hydrogen nucleon column density ($N_{\rm H}$) by multiplying by the mean extinction per H nucleon factor 5.8 $\times$$10^{21}$ cm-2/mag from Bohlin et al. (1978). Finally, we computed the hydrogen nucleon density ($n_{\rm H}$) by dividing by the width of each distance channel.

### v2

@@ 12. University of Vienna, Department of Astrophysics, Türkenschanzstraße 17, 1180 Wien, Austria
(Received 15 November 2024 / Accepted 4 February 2025)
We present a reconstruction of the line-of-sight motions of the local interstellar medium (ISM) based on the combination of a model of the three-dimensional dust density distribution within 1.25 kpc from the Sun and the Hi and CO line emission within Galactic latitudes $|b|$$\leq$ 5. We used the histogram of oriented gradient (HOG) method, a computer vision technique for evaluating the morphological correlation between images, to match the plane-of-the-sky dust distribution across distances with the atomic and molecular line emission. We identified a significant correlation between the 3D dust model and the line emission. We employed this correlation to assign line-of-sight velocities to the dust across density channels and produce a face-on map of the local ISM radial motions with respect to the local standard of rest (LSR). We find that most of the material in the 3D dust model follows the large-scale pattern of Galactic rotation; however, we also report local departures from the rotation pattern with standard deviations of 10.8 and 6.6 km s-1 for the Hi and CO line emission, respectively. The mean kinetic energy densities corresponding to these streaming motions are around 0.11 and 0.04 eV/cm3 from either gas tracer. Assuming homogeneity and isotropy in the velocity field, these values are within a factor of a few of the total kinetic energy density. These kinetic energy values are roughly comparable to other energy densities, thus confirming the near-equipartition in the local ISM. Yet, we identify energy and momentum overdensities of around a factor of ten concentrated in the Radcliffe Wave, the Split, and other local density structures. Although we do not find evidence of the local spiral arm’s impact on these energy overdensities, their distribution suggests the influence of large-scale effects that, in addition to supernova feedback, shape the energy distribution and dynamics in the solar neighborhood.
@@ 1 Introduction
Tracing the flows of matter and energy in the interstellar medium (ISM) is instrumental for understanding the structure and evolution of the Milky Way and other galaxies (see, for example, Ballesteros-Paredes et al., 2020; Tacconi et al., 2020; Chevance et al., 2023). Reconstructing the flow of gas in the ISM enables the identification of the mechanisms that make gas available for star formation and allows us to quantify the impact of galactic dynamics, magnetic fields, stellar winds, outflows, and supernovae (SNe) in the ISM (e.g., Field & Saslaw, 1965; Tomisaka, 1984; Koyama & Inutsuka, 2000; Blitz et al., 2007; Hennebelle et al., 2008; Klessen & Hennebelle, 2010; Dobbs et al., 2014; Krumholz et al., 2014; Schmidt et al., 2016; Hennebelle & Inutsuka, 2019). In this work, we present a reconstruction of the ISM line-of-sight (LOS) motions in the Solar neighborhood using state-of-the-art models of the three-dimensional dust density distribution, which we associate with the velocity information from the neutral atomic hydrogen (Hi) and carbon monoxide (CO) line emission using a machine vision tool to quantify the morphological similarities between distance and velocity channels (Soler et al., 2019).
Classic studies of the ISM distribution and kinematics rely on the assumption of circular motions around the Galactic center to compute what is usually called “kinematic distances” (see, for example, Oort et al., 1958; Roman-Duval et al., 2009; Sofue, 2011; Wenger et al., 2018; Hunter et al., 2024). Kinematic distances depend on the Galactic rotation model and are affected by the kinematic distance ambiguity (see, for example, Eilers et al., 2020; Reid, 2022). Moreover, the motion of the Galactic ISM is not purely rotational; non-circular streaming motions are induced by the Galactic bar, the spiral arms, and stellar feedback (Burton, 1971; Gómez, 2006; Moisés et al., 2011). The latter effect is expected to be prominent in the Solar neighborhood, where the motions introduced by supernovae (SNe)-blown bubbles and other local effects are potentially dominant over Galactic rotation as seen from the local standard of rest (LSR; see, for example, Zucker et al., 2023, and references therein).
@@
In this work, we associated the state-of-the-art 3D dust models presented by Edenhofer et al. (2024) with the LOS velocity information provided by Hi and CO surveys toward the Galactic plane. Following the pioneering work in Tchernyshyov & Peek (2017); Tchernyshyov et al. (2018), we adopted the term “kinetic tomography” to describe our reconstruction of the density and LOS velocity distribution. Our approach is based on the morphological similarity quantified by the histogram of oriented gradients (HOG) method (Soler et al., 2019). Using HOG, we linked the position-position-distance (PPD) 3D dust density cubes and the position-position-velocity (PPV) Hi and CO line emission cubes. The result is a position-position-distance-velocity (PPDV) hypercube, which we employ to characterize the local ISM motions along the line of sight. Line emission from ISM tracers and 3D dust have previously been combined to investigate the ISM in four dimensions (see, for example, Ivanova et al., 2021; Duchêne et al., 2023), but not with the level of reconstruction in the Edenhofer et al. (2024) maps, and never before with the HOG method.
@@
We used the Hi line emission at 21-cm wavelength, which traces both the cold, pre-molecular state before star formation and the warm, diffuse ISM before and after star formation (see, McClure-Griffiths et al., 2023, for a recent review). Since its discovery, Hi has been instrumental in studying the diffuse ISM in the Milky Way (Ewen & Purcell, 1951; Muller & Oort, 1951; Pawsey, 1951). Early observations of the Hi absorption against radio continuum sources revealed the presence of narrow, few-km s-1-wide, spectral features (e.g., Hagen et al., 1955; Clark, 1965). In emission, these narrow features appear on top of broader 10 to 20-km s-1-wide features (e.g., Matthews, 1957). These observations inspired the formulation of a “two phase” Hi model (Field et al., 1966; McKee & Ostriker, 1977), in which at the pressure of the ISM, the heating and cooling processes naturally lead to two thermally stable states: a dense cold neutral medium (CNM; $T$$\approx$ 50 K and $n$$\approx$ 50 cm-3) immersed in a diffuse warm neutral medium (WNM; $T$$\approx$ 8 000 K and $n$$\approx$ 0.3 cm-3).
@@
Part of our understanding of the Hi multiphase structure comes from studying absorption toward continuum sources. (see, for example, Strasser et al., 2007; Stanimirović et al., 2014; Murray et al., 2018). Comparisons between 21-cm Hi emission and absorption measurements indicate that, in the vicinity of the Sun, the WNM has roughly the same column density as the CNM (Falgarone & Lequeux, 1973; Liszt, 1983). Crucial additional information about the distribution of the CNM in and around MCs comes from the portion of the CNM sampled by the extended absorption of background Hi emission by cold foreground Hi, which is generically known as Hi self-absorption (HISA; Heeschen, 1955; Gibson et al., 2000; Seifried et al., 2020). Observations indicate that the HISA distribution is related to that of the CO emission, although the CNM it reveals corresponds to less than 15% than the total gas mass (see, for example, Gibson et al., 2005; Krčo et al., 2008; Wang et al., 2020; Syed et al., 2020).
@@
We also used the CO $(J$$=$$1$$\rightarrow$0)0)0 ) emission at 2.6 mm wavelength, which is the archetypal tracer of the cooler and denser ISM (Wilson et al., 1970). With its low excitation energy and critical density, CO provides an irreplaceable proxy for H2, which is the most abundant molecule in the Galaxy but is much harder to observe (Combes, 1991; Bolatto et al., 2013; Heyer & Dame, 2015). The focus on the molecular material traced by CO in star formation (SF) studies was cemented by the observed correlation between the SF and CO surface densities in nearby galaxies, which contrasts with the lack of correlation between SF and Hi at low surface densities in those objects (see, for example, Kennicutt, 1998; Wong & Blitz, 2002; Leroy et al., 2008; Schinnerer & Leroy, 2024).
@@
In this paper, we applied the HOG method to study the atomic and molecular gas motions toward the nearby Milky Way’s disk, defined as the Galactic latitude range $|b|$$\leq$ 5 within the 1.25-kpc extent of the Edenhofer et al. (2024) 3D dust reconstruction. The presentation of the data, analysis methods, and results is organized as follows. In Sec. 2, we present the 3D dust models and the Hi and CO observations used in the analysis. Section 3 summarizes the main aspects of the HOG method implementation for analyzing the Galactic plane. We describe the morphological correlation between the 3D dust models and the line emission in Sec. 4. In Sec. 5, we detail the streaming motions obtained with the HOG method and the energy and momentum densities derived from them. Section 6 discusses our results and their implications for understanding the ISM dynamics in the Solar neighborhood. We present our conclusions of this work in Sec. 7. We complement the main results of this work with the analysis shown in a set of appendices. Appendix A presents details on the HOG method’s error propagation and selection of parameters. We consider the effect of HISA in our results and the potential of the HOG method to identify these features in App. B. In App. C, we evaluate the effects of the fixed angular resolution and distance in the HOG results. Appendix D presents the HOG analysis of synthetic line emission and 3D density from a multiphase magnetohydrodynamic (MHD) MC simulation. Appendix E compares our results with the distances and line-of-sight velocities for the five maser sources within the studied volume. Finally, App. F presents the distance-velocity mapping for a few regions of interest, including areas toward the Galactic center and anticenter, where uncertainties in the kinematic distance estimates are considerable.
@@ 2 Data
In this paper, we use $v_{\rm LOS}$ to refer to the radial velocity inferred from the emission line Doppler shift relative to the LSR, usually identified as $v_{\rm LSR}$ in the Hi and CO observations. We introduce this convention to distinguish it from the general motion relative to the LSR, for which the three components of the velocity vector can be inferred in the Solar vicinity from the measurement of stellar motions (see, for example, Miret-Roig et al., 2022; Ratzenböck et al., 2023).
The selected $b$ range implies that we are studying the volume covered by the revolution of a trapezoid with base sizes 12.16 and 220.4 pc located at 69 and 1250 pc from the LSR, as illustrated in Fig. 2. The limits along the LOS are set by the boundaries of the Edenhofer et al. (2024) 3D dust model. At the furthest point in this analysis, we cover a region roughly within the characteristic CO scale height around the Solar circle ($\sim$ 100 pc, Heyer & Dame, 2015).
We consider the line emission in the range $-25$$<$$v_{\rm LOS}$$<$$25$ km s-1, which corresponds to the expected amplitude of LOS velocities for a Galactic rotation model for heliocentric distances within 1250 pc plus or minus a few kilometers per second. This restricted input range mitigates the spurious chance correlations produced by the limited angular resolution in the 3D dust reconstruction, as detailed in App A. However, it also introduces a generic limitation in our analysis, as some local ISM structures can have significant morphological correlations beyond the input $v_{\rm LOS}$ range. Thus, in practice, the restricted $v_{\rm LOS}$ input range limits the amplitude of reconstructed LOS motions for the local density structures. Until future higher-resolution 3D dust reconstructions break the ambiguities that produce chance correlations and expand the reliable input $v_{\rm LOS}$ range, our HOG-based reconstruction should be considered lower limits for the streaming motions and other derived quantities.
@@
The primary dataset enabling our study is the reconstruction of the 3D dust density between 69 and 1250 pc from the Sun presented in Edenhofer et al. (2024). This model uses the extinction estimates from Zhang et al. (2023b), which are primarily based on the Gaia satellite’s BP/RP spectra, with a spectral resolution ($R$) of roughly between 30 and 100 (Gaia Collaboration et al., 2023). Zhang et al. (2023b) forward-modeled the extinction, distance, and intrinsic parameters of each star given the combination of the Gaia spectra (Carrasco et al., 2021; De Angeli et al., 2023; Montegriffo et al., 2023) and infrared photometry from the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006), and unWISE, a processed catalog based on the observations from NASA’s Wide-field Infrared Survey Explorer (WISE) mission (Wright et al., 2010; Schlafly et al., 2019). The model was trained using a subset of stars observed with higher spectral resolution, $R$$\sim$ 1800, available from the Large Sky Area Multi-object Fibre Spectroscopic Telescope (LAMOST; Wang et al., 2022; Xiang et al., 2022).
@@
Edenhofer et al. (2024) obtains a spherical-coordinate reconstruction beyond 1 kpc while still resolving nearby dust clouds at parsec-scale resolution by implementing a new Gaussian process (GP) methodology to incorporate smoothness in a spherical coordinate system, mitigating fingers-of-god artifacts. The authors modeled the 3D distribution of differential extinction for stars in the Zhang et al. (2023b) catalog, assuming that the dust extinction distribution is spatially smooth. The posterior of their extinction model was reconstructed using variational inference and Gaussian variational inference (MGVI, Knollmüller & Enßlin, 2019; Leike & Enßlin, 2019). We refer to Edenhofer et al. (2024) for further data modeling and reconstruction details.
The result of the Edenhofer et al. (2024) model is a set of 12 samples drawn from the variational posterior of the 3D dust extinction distribution; that is, a set of 12 possible distributions consistent with the observations within the uncertainties. Each sample gives the value of the dust density in voxels arranged as a series of HEALPix [^1] HEALPix is the Hierarchical Equal Area isoLatitude Pixelization; see http://healpix.sf.net and Górski et al. (2005). spheres. There are 516 of these spheres, each corresponding to a distance logarithmically spaced grid between 69 to 1250 pc. Each sphere has a resolution parameter $N_{\rm side}$$=$ 256, corresponding to an angular size of around 13′7 for each voxel. We analyzed each of the 3D dust extinction samples, considering them independent realizations of 3D dust extinction, and reported the mean trends, as described in detail in App. A.
Using the tools in the Python healpy package, we produced a Cartesian projection of the 3D dust extinction covering the range $-180$$<$$l$$<$ 180 and $-5$$<$$b$$<$ 5 with a pixel size $\delta l$$=$$\delta b$$=$ 7.5. The reconstruction was made in terms of the unitless extinction defined by Zhang et al. (2023b), which we transformed into V-band extinction ($A_{\rm V}$) by multiplying with the 2.8 factor derived from their extinction curve. We then transformed $A_{\rm V}$ into hydrogen nucleon column density ($N_{\rm H}$) by multiplying by the mean extinction per H nucleon factor 5.8 $\times$$10^{21}$ cm-2/mag from Bohlin et al. (1978) and dividing by the total-to-selective extinction ratio, $R_{\rm V}$$\approx$ 3.1 (Savage & Mathis, 1979). Finally, we computed the hydrogen nucleon density ($n_{\rm H}$) by dividing by the width of each distance channel.